OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. M(0, 6), N(6, 0)

OpenStudy (anonymous):

Alright, so we have a line that passes through (0,6) and (6,0) So the slope of this line is -6/6, or just -1. In the slope-intercept equation of a line, we have y = -x + 6 To have it in the standard form of a line, we need x and y on the same side, so subtract x from both sides. y-x=6 There you go! The attached file is a picture of this line on a graph. :)

OpenStudy (anonymous):

Whoops, I messed up, it should be y+x = 6 because we add x to both sides, not subtract x from both sides.. Sorry about that.